OpenStudy (anonymous):

Which fraction is between 0 and 1/2? A.3/4 B.6/8 C.5/6 D.1/8

OpenStudy (anonymous):

Medal And Fan!

OpenStudy (chantysquirrel1129**):

Oh there ya go, now you got the right section! cx

OpenStudy (anonymous):

c';

OpenStudy (chantysquirrel1129**):

Make a number line! :3

OpenStudy (tkhunny):

It might help to note that 1/2 = 2/4 = 3/6 = 4/8

OpenStudy (anonymous):

Okaaaaayyyy then

OpenStudy (anonymous):

Hmm so

OpenStudy (chantysquirrel1129**):



OpenStudy (chantysquirrel1129**):

Let's start off with placing 1/2 on the number line c:

OpenStudy (anonymous):

oki

OpenStudy (chantysquirrel1129**):

Now, like what tk said, "1/2 = 2/4 = 3/6 = 4/8"

OpenStudy (anonymous):

Mhm

OpenStudy (chantysquirrel1129**):

The question is asking which fraction is BETWEEN 0 and 1/2! So if there are any answer choices that are equal to 1/2, cross them out :)

OpenStudy (chantysquirrel1129**):

So, let's start placing the fractions on the number line c:

OpenStudy (anonymous):

1/8?

OpenStudy (chantysquirrel1129**):

Are you asking if that's the answer? c:

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

If thats okay.

OpenStudy (anonymous):

?

OpenStudy (chantysquirrel1129**):

Well, ye, you are correct! :3

OpenStudy (chantysquirrel1129**):

Great work, peep! :D

OpenStudy (anonymous):

OOOOOHOHOHOHOHOHOOHHHOHOHOOHOHHHHH MUM GET THE CAMERAAAA

OpenStudy (anonymous):

Thanks sorry for that .0.

OpenStudy (anonymous):

Have any room for anymore questions?

OpenStudy (chantysquirrel1129**):

LOl what? xD It's ok! Keep up deh good work, dude C:

OpenStudy (chantysquirrel1129**):

Yes, just post another question on a NEW post!

OpenStudy (anonymous):

Which fraction is between 0 and ? A.4/6 B.1/3 C.7/10 D.3/5

OpenStudy (anonymous):

0 and 4/8

OpenStudy (anonymous):

Sorry

OpenStudy (chantysquirrel1129**):

On a new post lol cx